OpenStudy (anonymous):

Find the midpoint of the segment with the given endpoints. M(–2, 1) and O(–3, 2) A. (–2.5, 1.5) B. (–0.5, –0.5) C. (0.5, –0.5) D. (–3, –5)

OpenStudy (anonymous):

its D.

OpenStudy (anonymous):

K so a midpoint is basically an average of the x points and of the y points. The formula is x1+x2 divided by 2, y1+y2 divided by 2. If we do this we get- (-2+-3) (1+2) ------- & ----- 2 2 -5/2, 3/2= (-2.5,1.5)

OpenStudy (anonymous):

Ohhhh okay I get it! :D Okay can you help me with one more?

OpenStudy (anonymous):

I believe I know the answer but I need to know if I'm right.

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Wires are attached to a pole to make it more secure. The diagram shows one of those wires having a length of 220 feet. The angle of elevation from the ground to the top of the pole is 36. What is the height of the pole?

OpenStudy (anonymous):

A. 178.0 ft B. 374.3 ft C. 159.8 ft D. 129.3 ft

OpenStudy (anonymous):

This one is C what do you think?

OpenStudy (anonymous):

Sorry, but Idk this one :( I'm thinking maybe it has something to do w/ tan/cos/sin

OpenStudy (anonymous):

Hmmm okay ill figure it out tho... Thank you

OpenStudy (anonymous):

Sry :( I think though it has something to do with sine

OpenStudy (anonymous):

yeah I just put what I thought :) thanks tho...

OpenStudy (anonymous):

This can be solved with trigonometry. The hypotenuse has been given and the opposite side is needed. In this case, we will need to use sine. sin = opposite/hypotenuse let the opposite side length be y so sin(36) = y/220 220sin(36) = y plug that into a scientific calculator and you get 129.3ft so the answer is D